# Flat quasi-Frobenius Lie superalgebras

### v1

@@ 5.5.9. Odd flat double extensions of orthosympletic Lie superalgebras
In order to introduce the process of the odd flat double extension, we consider an odd linear map $\xi:\mathfrak{b}\rightarrow\mathfrak{b},$ and an even element $b_{0}\in\mathfrak{b}_{\bar{0}}$. For all $u\in{\mathfrak{b}}$, we define the linear map
(26) $D(u)=-(-1)^{|u|}(\xi^{*}+\xi)(u).$
@@
$(u+\alpha e)\diamond(v+\beta e):=u\star v+(-1)^{|v|}\omega(\xi(u),v)\,e,\qquad\forall\,u,v\in\mathfrak{b},\;\alpha,\beta\in\operatorname{\mathbb{K}},$
@@
(28) $\xi([u,v])=\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)),$
for all $u,v\in\mathfrak{b}$.
@@
(29) $\left\{\begin{array}[]{rclrcl}\xi([u,v])&=&\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)),&\left[\xi,\xi^{*}\right]&=&{\mathrm{R}}^{\star}_{b_{0}}-3\xi^{2},\\[5.69054pt] D(u\star v)&=&D(u)\star v+(-1)^{|u|}u\star D(v)+\xi(u)\star v,&{\mathrm{L}}^{\star}_{b_{0}}&=&-(\xi+\xi^{*})^{2},\\[5.69054pt] (2\xi+\xi^{*})(b_{0})&=&0,\end{array}\right$.
@@
$\xi([u,v])=\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)).$
@@
Furthermore, the natural symplectic product $\star$ on associated with $(\mathfrak{g},[\;,\;],\omega)$, is given by
@@
$\mathop{\mathrm{ad}}\nolimits_{u}^{k}(v)=(\mathop{\mathrm{ad}}\nolimits_{u}^{\mathfrak{b}})^{k}(v)+\big((-1)^{\alpha}\omega_{\mathfrak{b}}(\xi(u),\mathop{\mathrm{ad}}\nolimits_{u}^{\mathfrak{b}})^{k-1}(v))+(-1)^{|u|}\omega_{\mathfrak{b}}(\xi^{*}(u),\mathop{\mathrm{ad}}\nolimits_{u}^{\mathfrak{b}})^{k-1}(v))\big)e,$
@@ 5.5.16. Odd Flat double extension of flat periplectic Lie superalgebras
$(u+\alpha e)\diamond(v+\beta e):=u\star v+(-1)^{|v|}\omega(\xi(u),v)\,e,\qquad\forall\,u,v\in\mathfrak{b},\;\alpha,\beta\in\operatorname{\mathbb{K}},$
@@
(35) $\xi([u,v])=\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)),$
for all $u,v\in\mathfrak{b}$.
@@
(36) $\left\{\begin{array}[]{rclrcl}\xi([u,v])&=&\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)),&[\xi,\xi^{*}]&=&3\xi^{2}+{\mathrm{R}}^{\star}_{b_{0}},\\[5.69054pt] D(u\star v)&=&D(u)\star v+(-1)^{|u|}v\star D(v)+(-1)^{|u|}\xi(u)\star v,&{\mathrm{L}}^{\star}_{b_{0}}&=&(\xi+\xi^{*})^{2},\\[5.69054pt] 2\xi(b_{0})&=&\xi^{*}(b_{0}),\end{array}\right$.
@@
$\xi([u,v])=\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)).$
@@ 6. Flat quasi-Frobenius Lie superalgebras are nilpotent
With this observation in mind, let us prove our result by contradiction. According to Prop. 4.4.  Theorem, $(\mathfrak{g},[\;,\;])$ is solvable and $Z(\mathfrak{g})\cap[\mathfrak{g},\mathfrak{g}]=\{0\}$. We consider the sequence of vector subspaces defined by
@@
By virtue of Prop.  4.6.  Proposition, we have $C_{0}\neq\{0\}$, and according to Prop. 4.7.  Proposition, $C_{k}\subset N_{\ell}(\mathfrak{g},\star)$ for any $k\geq 1$.
@@
If $\omega$ is odd, then $\dim(\mathfrak{g}_{\bar{0}})=\dim(\mathfrak{g}_{\bar{1}})$. Assume $a\in Z(\mathfrak{g})_{\bar{0}}$. We can find an element $b\in\mathfrak{g}_{\bar{1}}$ such that $\omega(a,b)=1$ and $b,b]=\alpha a$, where $\alpha\in\mathbb{K}$. As $\omega$ is closed, we obtain
@@
which implies $\alpha=0$, and hence $(\mathfrak{g},[\;,\;])$ is abelian. The case where $a\in Z(\mathfrak{g})_{\bar{1}}$ is obvious.
[…]
@@ 2. $(ii)$
    $\pi_{\mathfrak{b}}(u),\pi_{\mathfrak{b}}(v)]_{\mathfrak{b}}=\pi_{\mathfrak{b}}([u,v])\quad\text{and}\quad\omega_{\mathfrak{b}}(\pi_{\mathfrak{b}}(u),\pi_{\mathfrak{b}}(v))=\omega(u,v),$
    for all $u,v\in I^{\perp}$.
@@
The Lie superalgebra $\mathfrak{b}$ obtained in Prop. 6.5.  Proposition is called the flat orthosymplectic quasi-Frobenius Lie superalgebra deduced from $I^{\perp}$ by means of $I$.
@@
$\omega_{\mathfrak{b}}(\xi(u),v)=\omega(u\star v,d)=\tfrac{1}{3}([u,v],d)+\tfrac{1}{3}\omega([u,d],v)=\tfrac{1}{3}\omega_{\mathfrak{b}}((\xi+\xi^{*}-D)(u),v),$
@@
The case, where $e$ is odd. Since $\omega$ is even and non-degenerate, there exists an element $d\in{\mathfrak{b}}_{\bar{1}}\setminus\{0\}$ such that $\omega(e,d)=\omega(d,e)=1.$
@@ 7.1. Four-dimensional flat quasi-Frobenius Lie superalgebras
Every flat orthosymplectic quasi-Frobenius Lie superalgebra of dimension three is abelian. Indeed, we suppose the Lie superalgbera is not abelian. Since $\omega$ is even, the even part $\mathfrak{g}_{0}$ has even dimension. It follows from Theorem 6.7.  Theorem that this $(\mathfrak{g},[\,,\,],\omega)$ is an even flat double extension of an one-dimensional flat orthosymplectic quasi-Frobenius Lie superalgebra $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ by means of $(\xi,b_{0})$. Since $\omega_{\mathfrak{b}}$ is even, it followsthe $\mathfrak{b}_{\bar{0}}=0$. Moreover, as both $\xi$ and $b_{0}$ are even, it follows that $\xi=0$ and $b_{0}=0$. Hence, from (21), all brackets vanish, and therefore $(\mathfrak{g},[\,,\,],\omega)$ is abelian, a contradiction.
@@
Let $(\mathfrak{b},[\,,\,]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ be the two-dimensional orthosymplectic quasi-Frobenius abelian Lie superalgebra.
@@
      $\xi=\begin{pmatrix}0&a\\ 0&0\end{pmatrix},\quad b_{0}=\alpha e_{1},\quad\omega=e_{1}^{*}\wedge e_{2}^{*},$
      where $\alpha\in\mathbb{K}$ and $a\neq 0$.
@@
    The pair $(\xi,b_{0})$ is odd admissible if and only if there exists a basis $\mathbb{B}=\{e_{1}\mid f_{1}\}$ of $\mathfrak{b}$ such that the following conditions are satisfied
    $\xi=0\quad b_{0}=\alpha e_{1}\quad\omega=e_{1}^{*}\wedge f_{1}^{*},\quad\text{where $\alpha\in\mathbb{K}$.}$
  • (iii)
@@
    $\begin{array}[]{lll}\text{(1)}&\xi=0,&b_{0}=\alpha\,e_{1},\quad\alpha\in\mathbb{K},\\[4.0pt] \text{(2)}&\xi=\begin{pmatrix}0&0\\[2.0pt] a&0\end{pmatrix},&b_{0}=0,\qquad a\neq 0,\\[10.0pt] \text{(3)}&\xi=\begin{pmatrix}0&a\\[2.0pt] 0&0\end{pmatrix},&b_{0}=\alpha\,e_{1},\qquad a\neq 0,\;\alpha\in\mathbb{K},\end{array}$
    where $\alpha\in\mathbb{K}$.
@@
$\omega=e_{1}^{*}\wedge f_{1}^{*},\qquad\xi(e_{1})=a\,e_{1},\qquad\xi(f_{1})=b\,f_{1},\quad b_{0}=\alpha e_{1}$
@@
Let $(\mathfrak{g},[\,,\,],\omega)$ be a flat quasi-Frobenius non-abelian Lie superalgebra of total dimension $4$. Theorems 6.7.  Theorem and 6.9.1.  Theorem, implies that such a superalgebra is an even or an odd flat double extension of the two-dimensional abelian quasi-Frobenius Lie superalgebra $(\mathfrak{b},[\,,\,]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ by means of $(\xi,b_{0})$. Following Prop. 7.1.3.  Proposition, we argue as follows.
@@
      Since $(\mathfrak{g},[\;,\;])$ is non-abelian, then $(\alpha,\beta)\neq(0,0)$. If for instance $\beta\neq 0$, then we choose
      $(x_{1},x_{2},x_{3},x_{4})=(-\frac{1}{\beta}e_{1},d,-e,-\beta e_{2}-\alpha e_{1}).$
      It follows that $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathbb{K}\oplus\operatorname{\mathcal{H}}_{3},[\;,\;],\omega)$.
@@
      which implies that $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathbb{K}\oplus\operatorname{\mathcal{H}}_{3},[\;,\;],\omega)$.
@@
      $(x_{1},x_{2}\mid y_{1},y_{2})=\left(-\frac{1}{a}d,2ae,-e_{2},-e_{1}\right).$
@@
      which implies that $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathfrak{g}^{2},[\;,\;],\omega)$.
@@
    $(x_{1},x_{2}\mid y_{1},y_{2})=(\frac{1}{\beta}e_{1},2\beta e_{2}+2\alpha e_{1},d,e).$
    Thus $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathfrak{g}^{2},[\;,\;],\omega)$.
@@
    $(x_{1},x_{2}\mid y_{1},y_{2})=(d,-\alpha e_{1},-\frac{1}{\alpha}f_{1},-e).$
    Thus, $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathfrak{g}^{3},[\;,\;],\omega)$.
@@
      $(x_{1},x_{2}\mid y_{1},y_{2})=(-2\alpha e_{1},e,d,\frac{1}{\alpha}f_{1},).$
      Thus $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathfrak{g}^{4},[\;,\;],\omega)$.
@@
      $(x_{1},x_{2}\mid y_{1},y_{2})=(e_{1},-ae,-\frac{1}{a}d,f_{1},).$
      Thus $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathfrak{g}^{3},[\;,\;],\omega)$.
@@
      $(x_{1},x_{2}\mid y_{1},y_{2})=(-\frac{2}{a}e,ae_{1},\frac{1}{a}f_{1},ad).$
@@
      $(x_{1},x_{2}\mid y_{1},y_{2})=\left(-\lambda e_{1}-\mu e,\;-\lambda\,e_{1}+\mu\,e,\;\frac{1}{2}\big(\lambda^{-1}\,d-\mu^{-1}\,f_{1}\big),\;\lambda^{-1}\,d+\mu^{-1}\,f_{1}\right),$
      where $\lambda=\frac{\sqrt[3]{\alpha^{2}}}{\sqrt[3]{a}}$ and $\mu=\frac{\sqrt[3]{a^{2}}}{\sqrt[3]{\alpha}}$. Thus, $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathfrak{g}^{4},[\;,\;],\omega)$.

### v2

@@ 5.5.9. Odd flat double extensions of orthosympletic Lie superalgebras
In order to introduce the process of the odd flat double extension, we consider an odd linear map $\xi:\mathfrak{b}\rightarrow\mathfrak{b},$ and an even element $b_{0}\in\mathfrak{b}_{\bar{0}}$. We define the linear map
(26) $D(u)=-(-1)^{|u|}(\xi^{*}+\xi)(u)\quad\text{For all $u\in{\mathfrak{b}}$.}$
@@
$(u+\alpha e)\diamond(v+\beta e):=u\star v+(-1)^{|v|}\omega(\xi(u),v)\,e,\qquad\text{ for all }u,v\in\mathfrak{b},\;\alpha,\beta\in\operatorname{\mathbb{K}},$
@@
(28) $\xi([u,v]_{\mathfrak{b}})=\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)),\quad\text{for all $u,v\in\mathfrak{b}$.}$
@@
(29) $\left\{\begin{array}[]{rclrcl}\xi([u,v]_{\mathfrak{b}})&=&\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)),&\left[\xi,\xi^{*}\right]&=&{\mathrm{R}}^{\star}_{b_{0}}-3\xi^{2},\\[5.69054pt] D(u\star v)&=&D(u)\star v+(-1)^{|u|}u\star D(v)+\xi(u)\star v,&{\mathrm{L}}^{\star}_{b_{0}}&=&-(\xi+\xi^{*})^{2},\\[5.69054pt] (2\xi+\xi^{*})(b_{0})&=&0,\end{array}\right$.
@@
$\xi([u,v]_{\mathfrak{b}})=\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)).$
@@
$\mathop{\mathrm{ad}}\nolimits_{u}^{k}(v)=(\mathop{\mathrm{ad}}\nolimits_{u}^{\mathfrak{b}})^{k}(v)+\big((-1)^{\alpha}\omega_{\mathfrak{b}}(\xi(u),(\mathop{\mathrm{ad}}\nolimits_{u}^{\mathfrak{b}})^{k-1}(v))+(-1)^{|u|}\omega_{\mathfrak{b}}(\xi^{*}(u),(\mathop{\mathrm{ad}}\nolimits_{u}^{\mathfrak{b}})^{k-1}(v))\big)e,$
@@ 5.5.16. Odd Flat double extension of flat periplectic Lie superalgebras
$(u+\alpha e)\diamond(v+\beta e):=u\star v+(-1)^{|v|}\omega(\xi(u),v)\,e,\qquad\text{for all }u,v\in\mathfrak{b},\;\alpha,\beta\in\operatorname{\mathbb{K}},$
@@
(35) $\xi([u,v]_{\mathfrak{b}})=\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)),\text{ for all $u,v\in\mathfrak{b}$.}$
@@
(36) $\left\{\begin{array}[]{rclrcl}\xi([u,v]_{\mathfrak{b}})&=&\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)),&[\xi,\xi^{*}]&=&3\xi^{2}+{\mathrm{R}}^{\star}_{b_{0}},\\[5.69054pt] D(u\star v)&=&D(u)\star v-(-1)^{|u|}u\star D(v)+(-1)^{|u|}\xi(u)\star v,&{\mathrm{L}}^{\star}_{b_{0}}&=&(\xi-\xi^{*})^{2},\\[5.69054pt] 2\xi(b_{0})&=&\xi^{*}(b_{0}),\end{array}\right$.
@@
$\xi([u,v]_{\mathfrak{b}})=\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)).$
@@ 6. Flat quasi-Frobenius Lie superalgebras are nilpotent
With this observation in mind, let us prove our result by contradiction. According to Prop. 4.5.  Theorem, $(\mathfrak{g},[\;,\;])$ is solvable and $Z(\mathfrak{g})\cap[\mathfrak{g},\mathfrak{g}]=\{0\}$. We consider the sequence of vector subspaces defined by
@@
By virtue of Prop.  4.8.  Proposition, we have $C_{0}\neq\{0\}$, and according to Prop. 4.9.  Proposition, $C_{k}\subset N_{\ell}(\mathfrak{g},\star)$ for any $k\geq 1$.
@@
If $\omega$ is odd, then $\dim(\mathfrak{g}_{\bar{0}})=\dim(\mathfrak{g}_{\bar{1}})$ by Lemma 2.1.  Lemma. Assume $a\in Z(\mathfrak{g})_{\bar{0}}$. We can find an element $b\in\mathfrak{g}_{\bar{1}}$ such that $\omega(a,b)=1$ and $b,b]=\alpha a$, where $\alpha\in\mathbb{K}$. As $\omega$ is closed, we obtain
@@
which implies $\alpha=0$, and therefore $b$ is also central. Hence, $(\mathfrak{g},[\;,\;])$ is abelian. In the case where $a\in Z(\mathfrak{g})_{\bar{1}}$, any element $b\in\mathfrak{g}_{\bar{0}}$ commutes with $a$ and is therefore central.
If $\omega$ is even, then $(\mathfrak{g},[\;,\;])$ is either purely even or purely odd, as $\omega_{|_{{\mathfrak{g}}_{\bar{0}}}}$ is also nondegenerate and therefore exists only if ${\mathfrak{g}}_{\bar{0}}$ is even-dimensional. In the case where $(\mathfrak{g},[\;,\;])$ is purely odd, the Lie superalgebra is abelian by definition. Let us assume that $(\mathfrak{g},[\;,\;])$ is purely even. Any 2-dimensional Lie algebra with a non-zero center must be abelian.
@@ 2. $(ii)$
    $\pi_{\mathfrak{b}}(u),\pi_{\mathfrak{b}}(v)]_{\mathfrak{b}}=\pi_{\mathfrak{b}}([u,v])\quad\text{and}\quad\omega_{\mathfrak{b}}(\pi_{\mathfrak{b}}(u),\pi_{\mathfrak{b}}(v))=\omega(u,v),\quad\text{ for all $u,v\in I^{\perp}$. }$
@@
The Lie superalgebra $\mathfrak{b}$ obtained in Prop. 6.5.  Proposition is called the flat quasi-Frobenius Lie superalgebra deduced from $I^{\perp}$ by means of $I$.
@@
$\omega_{\mathfrak{b}}(\xi(u),v)=\omega(u\star v,d)=\tfrac{1}{3}\omega([u,v],d)+\tfrac{1}{3}\omega([u,d],v)=\tfrac{1}{3}\omega_{\mathfrak{b}}((\xi+\xi^{*}-D)(u),v),$
@@
The case where $e$ is odd. Since $\omega$ is even and non-degenerate, there exists an element $d\in{\mathfrak{b}}_{\bar{1}}\setminus\{0\}$ such that $\omega(e,d)=\omega(d,e)=1.$
@@ 7.1. Four-dimensional flat quasi-Frobenius Lie superalgebras
Every flat orthosymplectic quasi-Frobenius Lie superalgebra of dimension three is abelian. Indeed, we suppose the Lie superalgbera is not abelian. Since $\omega$ is even, the even part $\mathfrak{g}_{0}$ has even dimension. It follows from Theorem 6.7.  Theorem that this $(\mathfrak{g},[\;,\;],\omega)$ is an even flat double extension of a one-dimensional flat orthosymplectic quasi-Frobenius Lie superalgebra $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ by means of $(\xi,b_{0})$. Since $\omega_{\mathfrak{b}}$ is even, it follows that $\mathfrak{b}_{\bar{0}}=0$. Moreover, as both $\xi$ and $b_{0}$ are even, it follows that $\xi=0$ and $b_{0}=0$. Hence, from (21), all brackets vanish, and therefore $(\mathfrak{g},[\,,\,],\omega)$ is abelian, a contradiction.
@@
Let $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ be a two-dimensional abelian orthosymplectic quasi-Frobenius Lie superalgebra.
@@
      $\xi=\begin{pmatrix}0&a\\ 0&0\end{pmatrix},\quad b_{0}=\alpha e_{1},\quad\omega=e_{1}^{*}\wedge e_{2}^{*},\quad\text{ where $\alpha\in\mathbb{K}$ and $a\neq 0$.}$
@@
    $\begin{array}[]{lll}\textup{(1)}&\xi=0,&b_{0}=\alpha\,e_{1},\quad\alpha\in\mathbb{K},\\[4.0pt] \textup{(2)}&\xi=\begin{pmatrix}0&0\\[2.0pt] a&0\end{pmatrix},&b_{0}=0,\qquad a\neq 0,\\[10.0pt] \textup{(3)}&\xi=\begin{pmatrix}0&a\\[2.0pt] 0&0\end{pmatrix},&b_{0}=\alpha\,e_{1},\qquad a\neq 0,\;\alpha\in\mathbb{K}.\end{array}$
@@
$\omega=e_{1}^{*}\wedge f_{1}^{*},\qquad\xi(e_{1})=a\,e_{1},\qquad\xi(f_{1})=b\,f_{1},\quad b_{0}=\alpha e_{1},$
@@
Let $(\mathfrak{g},[\;,\;],\omega)$ be a flat quasi-Frobenius non-abelian Lie superalgebra of total dimension $4$. Theorems 6.7.  Theorem and 6.9.1.  Theorem, implies that such a superalgebra is an even or an odd flat double extension of the two-dimensional abelian quasi-Frobenius Lie superalgebra $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ by means of $(\xi,b_{0})$. Following Props. 7.1.3.  Proposition and 7.1.4.  Proposition, we argue as follows.
@@
      Since $(\mathfrak{g},[\;,\;])$ is non-abelian, then $\left(\alpha,\beta\right)\neq(0,0)$. If for instance $\beta\neq 0$, then we choose
      $(x_{1},x_{2},x_{3},x_{4})=\left(-\frac{1}{\beta}e_{1},d,-e,-\beta e_{2}-\alpha e_{1}\right).$
      It follows that $(\mathfrak{g},[\;,\;],\omega)$ is symplectomorphic to $(\mathbb{K}\oplus\operatorname{\mathcal{H}}_{3},[\;,\;],\omega)$.
@@
      which implies that $(\mathfrak{g},[\;,\;],\omega)$ is symplectomorphic to $(\mathbb{K}\oplus\operatorname{\mathcal{H}}_{3},[\;,\;],\omega)$.
@@
      $(x_{1},x_{2},y_{1},y_{2})=\left(-\frac{1}{a}d,2ae,-e_{2},-e_{1}\right).$
@@
      which implies that $(\mathfrak{g},[\;,\;],\omega)$ is symplectomorphic to $(\mathfrak{g}^{2},[\;,\;],\omega)$.
@@
    $(x_{1},x_{2},y_{1},y_{2})=\left(\frac{1}{\beta}e_{1},2\beta e_{2}+2\alpha e_{1},d,e\right).$
    Thus, $(\mathfrak{g},[\;,\;],\omega)$ is symplectomorphic to $(\mathfrak{g}^{2},[\;,\;],\omega)$.
@@
    $(x_{1},x_{2},y_{1},y_{2})=\left(d,-\alpha e_{1},-\frac{1}{\alpha}f_{1},-e\right).$
    Thus, $(\mathfrak{g},[\;,\;],\omega)$ is symplectomorphic to $(\mathfrak{g}^{3},[\;,\;],\omega)$.
@@
      $(x_{1},x_{2},y_{1},y_{2})=\left(-2\alpha e_{1},e,d,\frac{1}{\alpha}f_{1},\right).$
      Thus, $(\mathfrak{g},[\;,\;],\omega)$ is symplectomorphic to $(\mathfrak{g}^{4},[\;,\;],\omega)$.
@@
      $(x_{1},x_{2},y_{1},y_{2})=\left(e_{1},-ae,-\frac{1}{a}d,f_{1},\right).$
      Thus, $(\mathfrak{g},[\;,\;],\omega)$ is symplectomorphic to $(\mathfrak{g}^{3},[\;,\;],\omega)$.
@@
      $(x_{1},x_{2},y_{1},y_{2})=\left(-\frac{2}{a}e,ae_{1},\frac{1}{a}f_{1},ad\right).$
@@
      $(x_{1},x_{2},y_{1},y_{2})=\left(-\lambda e_{1}-\mu e,\;-\lambda\,e_{1}+\mu\,e,\;\frac{1}{2}\big(\lambda^{-1}\,d-\mu^{-1}\,f_{1}\big),\;\lambda^{-1}\,d+\mu^{-1}\,f_{1}\right),$
      where $\lambda=\frac{\sqrt[3]{\alpha^{2}}}{\sqrt[3]{a}}$ and $\mu=\frac{\sqrt[3]{a^{2}}}{\sqrt[3]{\alpha}}$.
      Thus, $(\mathfrak{g},[\;,\;],\omega)$ is symplectomorphic to $(\mathfrak{g}^{4},[\;,\;],\omega)$.